# Waveform Optimization and Beam Focusing for Near-field Wireless Power Transfer with Dynamic Metasurface Antennas and Non-linear Energy Harvesters

### v2

@@ II-A Transmit Antenna Architectures
The transmitter is equipped with a uniform planar array (UPA) and $N_{rf}\geq M$ RF chains. The radiating elements are spaced uniformly, with $N_{h}$ and $N_{v}$ being the number of elements in the horizontal and vertical direction, respectively. Thus, the total number of elements is $N=N_{v}\times N_{h}$. Two types of transmit antenna architectures are considered: 1) Fully-digital architecture, which requires a dedicated RF chain for each radiating element, thus $N_{rf}=N$, as shown in Fig. 2a. In a fully-digital architecture, there is a single-stage beamforming process. Herein, $\omega_{i,n}$ is the complex weight of the $n$th frequency tone of the $i$th muti-tone waveform. Despite the high deployment cost and complexity, a fully-digital structure offers the highest number of degrees of freedom in beamforming.
Figure 2: Transmit antenna architectures. (a) fully-digital architecture (left) (b) DMA-assisted architecture (right).
2) DMA-assisted architecture, which comprises $N_{v}$ waveguides, each fed by a dedicated RF chain and composed of $N_{h}$ configurable metamaterial elements. Therefore, the number of RF chains, and consequently the cost and complexity, is considerably reduced compared to digital structures, making DMA suitable for mMIMO applications. Notice that DMA-assisted systems employ a two-stage beamforming process, i.e., digital beamforming, followed by the tuning of the amplitude/phase variations introduced by the metamaterial elements, as illustrated in Fig. 2b. Herein, $q_{i,l}$ is the tunable frequency response of the $l$th metamaterial element in the $i$th waveguide, while $h_{i,l}$ is the corresponding waveguide propagation loss, which will be explained in detail later.
@@ II-B Channel Model
[…]
Notice that one of the main applications of WPT systems is in indoor environments with line-of-sight (LOS) and near-field communication, e.g., restaurants, warehouses, and shopping malls. Thus, we employ the near-field LOS channel model described in [27]. The Cartesian coordinate of the $l$th radiating element in the $i$th row is $\mathbf{g}_{i,l}=[x_{i,l},y_{i,l},z_{i,l}]^{T}$. Additionally, $i=1,2,\ldots,N_{v}$ and $l=1,2,\ldots,N_{h}$. The channel coefficient between user $m$ and the $l$th element in the $i$th row at the $n$th sub-carrier is given by
@@
where ${\frac{2\pi}{\lambda_{n}}d_{i,l,m}}$ is the phase shift caused by the propagation distance of the $n$th tone, with wavelength $\lambda_{n}$, and $d_{i,l,m}=|\mathbf{g}_{m}-\mathbf{g}_{i,l}|$ is the Euclidean distance between the element and the user located at $\mathbf{g}_{m}$. Moreover,
@@
$G_{t}=2(b+1)$ is the transmit antenna gain, and $b$ denotes the boresight gain, which depends on the specifications of the antenna elements. Note that the channel coefficient becomes $A_{m}e^{-j\psi_{i,l,m}}$ for far-field communication, where $A_{m}$ only depends on the distance of the user $m$ from the transmitter and $\psi_{i,l,m}$ is solely determined by the user direction and the relative disposition of the antenna elements within the array.
@@ II-C Transmit & Receive Signals
$x_{i}(t)=\sum_{n=1}^{N_{f}}{\omega_{i,n}e^{j2\pi f_{n}t}},\quad i=1,\ldots,N_{rf}.$ (5)
The HPA introduces signal distortion and models such as the Rapp model [29] capture this non-linearity. It is shown in [21] that when the HPA operates in the non-linear regime, choosing a single-carrier waveform is preferred to a multi-carrier one. The reason is that a single-carrier waveform is less deteriorated by the adverse effect of the signal distortion caused by the HPA when operating in the non-linear regime. On the other hand, when HPAs operate in the linear regime, thus, not causing amplitude and phase distortion in the signal, multi-carrier waveforms are preferred since they leverage the rectifier’s non-linearity and enhance the harvested power performance.
@@ II-C1 Fully-Digital Architecture
where ${u{(i,l)}}=(i-1)N_{h}+l$. Thereby, the RF signal at the $m$th receiver when exploiting the fully-digital architecture can be expressed as
@@ II-C2 DMA-assisted Architecture
$q_{i,l}\in\mathcal{Q}=\Big{\{}{(j+e^{j\phi_{i,l}})}/{2}\Big{|}\phi_{i,l}\in[0,2\pi]\Big{\}},\quad\forall i,l.$ (9)
@@
Herein, microstrip lines are used as waveguides, similar to [31, 27]. The propagation model of the signal within a microstrip is expressed as $h_{i,l}=e^{-(l-1)d_{l}(\alpha_{i}+j\beta_{i})}$, where $d_{l}$ is the inter-element distance, $\alpha_{i}$ represents the waveguide attenuation coefficient, and $\beta_{i}$ is the propagation constant. The mathematical model of the DMA is represented in Fig. 2b.
@@
$x_{i,l}^{DMA}(t)$ $\!=\!G\Re\left\{h_{i,l}q_{i,l}x_{i}^{hpa}(t)\right\}\!=\!G\sum_{n=1}^{N_{f}}\Re\left\{h_{i,l}q_{i,l}{\omega_{i,n}e^{j2\pi f_{n}t}}\right\}.$ (10)
Furthermore, the RF signal received at the $m$th user in the DMA-assisted system is given by
@@ II-D Rectenna
$v_{{o},{m}}=\sum_{i\ even,i\geq 2}^{n_{0}}K_{i}\mathbb{E}\bigl{\{}y_{m}(t)^{i}\bigr{\}},$ (14)
@@
$v_{{o},{m}}=K_{2}\mathbb{E}\bigl{\{}y_{m}(t)^{2}\bigr{\}}+K_{4}\mathbb{E}\bigl{\{}y_{m}(t)^{4}\bigr{\}}$ (15)
and the DC power at the $m$th receiver is given by
@@
where $R_{L}$ is the load impedance of the rectifier, while $y_{m}(t)$ is equal to $y_{m}^{DMA}(t)$ and $y_{m}^{FD}(t)$ in the DMA and fully-digital.
@@ II-E Power Consumption Model
Lowering the system power consumption is desirable for increasing the end-to-end PTE and this greatly depends on the power consumption of the HPA. Let us denote the maximum efficiency and the maximum output power of a class-B HPA as $\bar{\eta}$ and $P_{max}$, respectively. Hereby, the efficiency of the $i$th HPA at time $t$ is expressed as $\eta_{i}(t)={\bar{\eta}}\sqrt{{P_{out,i}(t)}/{P_{max}}}$ [32], where $P_{out,i}(t)$ is the output power of the HPA. Then, the corresponding power consumption is
@@
Notice that each HPA is in charge of supplying the amount of radiating power by all of the antenna elements, which are fed by its corresponding RF chain. Thus, $P_{out,i}(t)$ is $\sum_{l=1}^{N_{h}}|x^{DMA}_{i,l}(t)|^{2}$ in DMA-assisted architecture, and $|x^{FD}_{i,l}(t)|^{2}$ denotes the output power of the RF chain connected to the $l$th element in the $i$th row.
@@
$P_{c}=\sum_{i=1}^{N_{rf}}\mathbb{E}\bigl{\{}P_{{hpa},i}(t)\bigr{\}}+P_{in},$ (18)
@@ III-A Problem Formulation
The goal is to obtain a minimum-power waveform and beamforming design to meet the EH requirements of the users. Thus, by utilizing (16) and substituting (17) in (18), the optimization problem can be formulated as
$\operatorname*{minimize}_{\mathcal{V}}\quad$ $\frac{\sqrt{P_{max}}}{\bar{\eta}}\sum_{i=1}^{N_{rf}}\mathbb{E}\bigl{\{}\sqrt{P_{out,i}}(t)\bigr{\}}+\sum_{i=1}^{N_{rf}}\sum_{n=1}^{N_{f}}|\omega_{i,n}|^{2}$ (19a)
@@
where $\mathcal{V}$ is the set of optimization variables, which is equal to $\{\omega_{i,n},q_{i,l}\}_{\forall i,l,n}$ and $\{\omega_{i,n}\}_{\forall i,n}$ for DMA-assisted and fully-digital architectures, respectively. Problem (III-A) deals with extensive non-linearity since the objective and constraints are highly non-linear and non-convex functions due to the signal model, rectenna non-linearity, and the coupling between the digital and analog beamforming variables in DMA.
@@ III-B1 Optimization with fixed $\{q_{i,l}\}_{\forall i,l}$
$\operatorname*{minimize}_{\{\omega_{i,n}\}_{\forall i,n}}\quad$ $\frac{\sqrt{P_{max}}}{\bar{\eta}}\sum_{i=1}^{N_{v}}\mathbb{E}\biggl{\{}\sqrt{\sum_{l=1}^{N_{h}}x^{DMA}_{i,l}(t)^{2}}\biggr{\}}\!+\!\sum_{i=1}^{N_{v}}\sum_{n=1}^{N_{f}}|\omega_{i,n}|^{2}$ (20a)
subject to $\bigl{(}{K_{2}\mathbb{E}\bigl{\{}y_{m}(t)^{2}\bigr{\}}+K_{4}\mathbb{E}\bigl{\{}y_{m}(t)^{4}\bigr{\}}\bigr{)}}^{2}\geq{R_{L}}\bar{P}_{m},\forall m.$ (20b)
The problem is still highly non-linear and non-convex. Interestingly, it can be easily verified that $v_{o,m}$ is convex and increasing with respect to ${y_{m}(t)}^{2}$, and ${y_{m}(t)}$ is affine with respect to $\bar{\mathbf{w}}$. Therefore, $v_{o,m}$ is a convex function with respect to $\bar{\mathbf{w}}$ given a fixed configuration for metamaterial elements [21]. Thus, although (20b) is a non-convex constraint, its left-hand side consists of a convex function. These properties motivate us to adapt the SCA method [33] to optimize the problem iteratively. Specifically, the objective and constraints can be approximated by their Taylor expansion.
Theorem 1.
By defining $S_{i}=\sum_{l=1}^{N_{h}}\mathbb{E}\bigl{\{}x^{DMA}_{i,l}(t)^{2}\bigr{\}}$, we can write
$\sum_{i=1}^{N_{rf}}\mathbb{E}\biggl{\{}\sqrt{\sum_{l=1}^{N_{h}}x^{DMA}_{i,l}(t)^{2}}\biggr{\}}\leq\sum_{i=1}^{N_{rf}}\tilde{f}\bigl{(}S_{i},S_{i}^{(0)}),$ (21)
where
$\tilde{f}\bigl{(}S_{i},S_{i}^{(0)})=\sqrt{S_{i}^{(0)}}+\frac{1}{2\sqrt{S_{i}^{(0)}}}\bigl{(}S_{i}-S_{i}^{(0)}\bigr{)},\forall i\vspace{-2mm}$ (22)
is the first order Taylor expansion of ${\sqrt{S_{i}}}$ at point $S_{i}^{(0)}$.
Proof.
The inequality can be proved using Jensen’s inequality [34] and the fact that ${\sqrt{S_{i}}}$ is a concave function with respect to ${S_{i}}$, while the first-order Taylor expansion of a function is greater than or equal to the function value at each point. ∎
Then, we rewrite $v_{o,m}$ considering fixed $q_{i,l}$. For this, we define $\mathbf{a}_{m,n}=\boldsymbol{\gamma}_{m,n}\odot\mathbf{q}\odot\mathbf{h}$ and leverage the fact that the average power of a signal is equal to the power of its spectrum. Thus, we have [6]
$\mathbb{E}\bigl{\{}y_{m}(t)^{2}\bigr{\}}=g_{m,1}(\{\bar{\mathbf{w}}_{n}\}_{\forall n})=\frac{G^{2}}{2}\sum_{n}|\mathbf{a}^{T}_{m,n}\bar{\mathbf{w}}_{n}|^{2},$ (23)
$\mathbb{E}\bigl{\{}y_{m}(t)^{4}\bigr{\}}=g_{m,2}(\{\bar{\mathbf{w}}_{n}\}_{\forall n})=\frac{3G^{4}}{8}\sum_{\begin{subarray}{c}n_{0},n_{1},n_{2},n_{3}\\ n_{0}+n_{1}=n_{2}+n_{3}\end{subarray}}...$ (26)
@@
Similarly, it can be verified that
$\mathbb{E}\bigl{\{}x^{DMA}_{i,l}(t)^{2}\bigr{\}}=\frac{G^{2}}{2}\sum_{n}|q_{i,l}h_{i,l}w_{i,n}|^{2}.$ (30)
@@
$\operatorname*{minimize}_{\begin{subarray}{c}\{\bar{\mathbf{w}}_{n}\}_{\forall n}\\ \{{v}_{o,m}\}_{\forall m}\\ \{S_{i}\}_{\forall i}\end{subarray}}\quad$ $\frac{\sqrt{P_{max}}}{\bar{\eta}}\sum_{i=1}^{N_{rf}}\tilde{f}\bigl{(}S_{i},S_{i}^{(0)})+\sum_{i=1}^{N_{rf}}\sum_{n=1}^{N_{f}}|\omega_{i,n}|^{2}$ (31d)
@@
$\sum_{n}\tilde{g}_{m,n}(\bar{\mathbf{w}}_{n}^{(0)})\bigl{(}\bar{\mathbf{w}}_{n}-\bar{\mathbf{w}}_{n}^{(0)}\bigr{)},\quad\forall m$ (31e)
$\frac{G^{2}}{2}\sum_{n=1}^{N_{f}}\sum_{l=1}^{N_{h}}\bigl{|}q_{i,l}h_{i,l}\bar{\mathbf{w}}_{n}[i]\bigr{|}^{2}\leq S_{i},i=1,\ldots,N_{v},$ (31f)
$\bar{\mathbf{w}}_{n}[(i-1)N_{h}+l]=\bar{\mathbf{w}}_{n}[i],\forall i,l=1,\ldots,N_{h}.$ (31g)
Notice that by utilizing (21) and the fact that (20a) consists of two positive terms, one can verify that (31d) serves as an upper bound for (20a), thus, minimizing the latter leads to minimizing the former. Moreover, the inequality in (28) ensures that the solution to this problem is a feasible solution of (III-B1) at each point. Interestingly, the problem [^2] Note that for the sake of notation and facilitating the reader’s understanding, we kept the problem in the vector form and introduced the constraint (31g) into the problem. However, the problem can be easily converted to scalar form, which removes the mentioned constraint. has become convex and can be solved at a given point by standard convex optimization tools, e.g., CVX [35]. Moreover, the solution can be iteratively updated using the SCA algorithm [33].

### v3

@@ II-A Transmit Antenna Architectures
The transmitter is equipped with a uniform planar array (UPA) and $N_{rf}\geq M$ RF chains. The radiating elements are spaced uniformly, with $N_{h}$ and $N_{v}$ being the number of elements in the horizontal and vertical direction, respectively. Thus, the total number of elements is $N=N_{v}\times N_{h}$. Two types of transmit antenna architectures are considered:
1) Fully-digital architecture, which requires a dedicated RF chain for each radiating element, thus $N_{rf}=N$, as shown in Fig. 2a. In a fully-digital architecture, there is a single-stage beamforming process. Despite the high deployment cost and complexity, a fully-digital structure offers the highest number of degrees of freedom in beamforming.
2) DMA-assisted architecture, which comprises $N_{v}$ waveguides, each fed by a dedicated RF chain and composed of $N_{h}$ configurable metamaterial elements. Therefore, the number of RF chains, and consequently the cost and complexity, is considerably reduced compared to digital structures, making DMA suitable for massive MIMO applications. Notice that DMA-assisted systems employ a two-stage beamforming process, i.e., digital beamforming, followed by the tuning of the amplitude/phase variations introduced by the metamaterial elements, as illustrated in Fig. 2b. Herein, $q_{i,l}$ is the tunable frequency response of the $l$th metamaterial element in the $i$th waveguide, while $h_{i,l}$ is the corresponding waveguide propagation loss, which will be explained in detail later.
@@ II-B Channel Model
In wireless communications, the region where the receivers are located between the Fraunhofer and Fresnel distances denoted respectively as $d_{fr}$ and $d_{fs}$, is the radiative near-field region, which is referred to as the near-field region in the following. Specifically, a receiver at distance $d$ from a transmitter experiences near-field conditions if
$3]{\frac{D^{4}}{8\lambda_{1}}}=d_{fs}<d<d_{fr}=\frac{2D^{2}}{\lambda_{1}},$
where $D$ is the antenna diameter, i.e., the largest size of the antenna aperture, $\lambda_{1}=c/{f_{1}}$ is the corresponding wavelength to the system operating frequency, and $c$ is the speed of light. Notice that both system frequency and antenna form factor influence the region of operation. Therefore, by moving toward higher frequencies, e.g., millimeter wave and sub-THz bands, and/or utilizing larger antenna arrays, the far-field communication assumption regarding planar wavefronts may not be valid anymore. Instead, wavefronts impinging a receive node may be strictly spherical, thus, with advanced capabilities to focus the transmit signals on specific spatial points rather on spatial directions. Thus, we use the term beam focusing since we consider the system to operate in the radiative near-field region.
Notice that one of the main applications of WPT systems is in indoor environments with line-of-sight and near-field communication, e.g., restaurants, warehouses, and shopping malls. Thus, we employ the near-field line-of-sight channel model described in [30]. The Cartesian coordinate of the $l$th radiating element in the $i$th row is $\mathbf{g}_{i,l}=[x_{i,l},y_{i,l},z_{i,l}]^{T}$. Additionally, $i=1,2,\ldots,N_{v}$ and $l=1,2,\ldots,N_{h}$. The channel coefficient between ER $m$ and the $l$th element in the $i$th row at the $n$th sub-carrier is given by
@@
where ${\frac{2\pi}{\lambda_{n}}d_{i,l,m}}$ is the phase shift caused by the propagation distance of the $n$th tone, with wavelength $\lambda_{n}$, and $d_{i,l,m}=|\mathbf{g}_{m}-\mathbf{g}_{i,l}|$ is the Euclidean distance between the element and the ER located at $\mathbf{g}_{m}$. Moreover,
@@
$G_{t}=2(b+1)$ is the transmit antenna gain, and $b$ denotes the boresight gain, which depends on the specifications of the antenna elements. Note that the channel coefficient becomes $A_{m}e^{-j\psi_{i,l,m}}$ for far-field communication, where $A_{m}$ only depends on the distance of the receiver $m$ from the transmitter and $\psi_{i,l,m}$ is solely determined by the ER direction and the relative disposition of the antenna elements within the array. Notably, the radiation profile might be more complicated in practice, e.g., due to back lobes. However, this work considers perfect CSI at the transmitter side; thus, the proposed frameworks apply also to other radiation profiles and channel models and we adapt a basic ideal radiation profile without loss of generality.
@@ II-C Transmit & Receive Signals
$x_{i}(t)=\sum_{n=1}^{N_{f}}{\omega_{i,n}e^{j2\pi f_{n}t}},\quad i=1,\ldots,N_{rf},$ (5)
where $\omega_{i,n}$ is the complex weight of the $n$th frequency tone of the $i$th muti-tone waveform. The HPA introduces signal distortion and models such as the Rapp model [32] capture this non-linearity. It is shown in [24] that when the HPA operates in the non-linear regime, choosing a single-carrier is preferred to a multi-carrier waveform. The reason is that a single-carrier waveform is less deteriorated by the adverse effect of the signal distortion caused by the HPA when operating in the non-linear regime. On the other hand, when HPAs operate in the linear regime, thus, not causing amplitude and phase distortion to the signal, multi-carrier waveforms are preferred since they leverage the rectifier’s non-linearity and enhance the harvested power performance.
@@ II-C1 Fully-Digital Architecture
where ${u{(i,l)}}=(i-1)N_{h}+l$. Thereby, the RF signal at the $m$th ER when exploiting the fully-digital architecture can be expressed as
@@ II-C2 DMA-assisted Architecture
$q_{i,l}\in\mathcal{Q}=\Big{\{}{(j+e^{j\phi_{i,l}})}/{2}\ \Big{|}\ \phi_{i,l}\in[0,2\pi]\Big{\}},\quad\forall i,l,$ (9)
@@
Herein, microstrip lines are used as waveguides, similar to [34, 30]. The propagation model of the signal within a microstrip is expressed as $h_{i,l}=e^{-(l-1)d_{l}(\alpha_{i}+j\beta_{i})}$, where $d_{l}$ is the inter-element distance, $\alpha_{i}$ represents the waveguide attenuation coefficient, and $\beta_{i}$ is the propagation constant. The mathematical model of the DMA is represented in Fig. 2b. For the sake of simplicity and since the goal is to investigate the power consumption of DMAs for RF-WPT purposes, we consider in this work a simplified case with no mutual coupling between the elements of the metasurface. However, when investigating DMAs from an electromagnetic perspective, one should carefully consider the impact of mutual coupling [35].
@@
$x_{i,l}^{DMA}(t)$ $=G\Re\left\{h_{i,l}q_{i,l}x_{i}^{hpa}(t)\right\}$
$=G\sum_{n=1}^{N_{f}}\Re\left\{h_{i,l}q_{i,l}{\omega_{i,n}e^{j2\pi f_{n}t}}\right\}.$ (10)
Furthermore, the RF signal received at the $m$th ER in the DMA-assisted system is given by
@@ II-D Rectenna
$v_{{o},{m}}=\sum_{i\ even,i\geq 2}^{n_{0}}K_{i}\mathbb{E}_{t}\bigl{\{}y_{m}(t)^{i}\bigr{\}},$ (14)
@@
$v_{{o},{m}}=K_{2}\mathbb{E}_{t}\bigl{\{}y_{m}(t)^{2}\bigr{\}}+K_{4}\mathbb{E}_{t}\bigl{\{}y_{m}(t)^{4}\bigr{\}}$ (15)
and the DC power at the $m$th ER is given by
@@
where $R_{L}$ is the load impedance of the rectifier, while $y_{m}(t)$ is equal to $y_{m}^{DMA}(t)$ and $y_{m}^{FD}(t)$ in the DMA-assisted and fully-digital architectures, respectively.
@@ II-E Power Consumption Model
Lowering the system power consumption is desirable for increasing the end-to-end PTE and this greatly depends on the HPA design and operation. Let us denote the maximum efficiency and the maximum output power of a class-B HPA as $\bar{\eta}$ and $P_{max}$, respectively. Hereby, the efficiency of the $i$th HPA at time $t$ is expressed as $\eta_{i}(t)={\bar{\eta}}\sqrt{{P_{out,i}(t)}/{P_{max}}}$ [36], where $P_{out,i}(t)$ is the output power of the HPA. Then, the corresponding power consumption is
@@
Notice that each HPA is in charge of supplying the radiating power by all the antenna elements, which are fed by its corresponding RF chain. Thus, $P_{out,i}(t)$ is $\sum_{l=1}^{N_{h}}{x^{DMA}_{i,l}(t)}^{2}$ in DMA-assisted architecture, and ${x^{FD}_{i,l}(t)}^{2}$ denotes the output power of the RF chain connected to the $l$th element in the $i$th row.
@@
$P_{c}=\sum_{i=1}^{N_{rf}}\mathbb{E}_{t}\bigl{\{}P_{{hpa},i}(t)\bigr{\}}+P_{in},$ (18)
@@ III-A Problem Formulation
The goal is to obtain a minimum-power waveform and beamforming design to meet the EH requirements of the ERs. For this, we assume that the locations of the receivers are fixed and known at the transmitter, which leads to the perfect CSI assumption. Thus, by utilizing (16) and substituting (17) into (18), the optimization problem can be formulated as
$\operatorname*{minimize}_{\mathcal{V}}\quad$ $\frac{\sqrt{P_{max}}}{\bar{\eta}}\sum_{i=1}^{N_{rf}}\mathbb{E}_{t}\bigl{\{}\sqrt{P_{out,i}(t)}\bigr{\}}+\sum_{i=1}^{N_{rf}}\sum_{n=1}^{N_{f}}|\omega_{i,n}|^{2}$ (19a)
@@
where $\mathcal{V}$ is the set of optimization variables, which is equal to $\{\omega_{i,n},q_{i,l}\}_{\forall i,l,n}$ and $\{\omega_{i,n}\}_{\forall i,n}$ for DMA-assisted and fully-digital architectures, respectively. Problem (19) deals with extensive non-linearity since the objective and constraints are highly non-linear and non-convex functions due to the signal model, rectenna non-linearity, and the coupling between the digital and analog beamforming variables in DMA. Notice that this problem formulation is equivalent to maximizing the PTE. Specifically, the solution will force the DC harvested power of the ERs to be equal to the EH requirement ($\bar{P}_{m}$). Thus, the numerator of (1) is constant, and decreasing the power consumption leads to maximizing the PTE.
@@ III-B1 Optimization with fixed $\{q_{i,l}\}_{\forall i,l}$
$\operatorname*{minimize}_{\{\omega_{i,n}\}_{\forall i,n}}\quad$ $\frac{\sqrt{P_{max}}}{\bar{\eta}}\sum_{i=1}^{N_{v}}\mathbb{E}_{t}\biggl{\{}\sqrt{\sum_{l=1}^{N_{h}}x^{DMA}_{i,l}(t)^{2}}\biggr{\}}\!+\!\sum_{i=1}^{N_{v}}\sum_{n=1}^{N_{f}}|\omega_{i,n}|^{2}$ (20a)
subject to ${K_{2}\mathbb{E}_{t}\bigl{\{}y_{m}(t)^{2}\bigr{\}}+K_{4}\mathbb{E}_{t}\bigl{\{}y_{m}(t)^{4}\bigr{\}}}\geq\sqrt{{R_{L}}\bar{P}_{m}},\forall m.$ (20b)
The problem is still highly non-linear and non-convex. Interestingly, it can be easily verified that $v_{o,m}$ is convex and increasing with respect to ${y_{m}(t)}^{2}$, and ${y_{m}(t)}$ is affine with respect to $\bar{\mathbf{w}}$. Therefore, $v_{o,m}$ is a convex function with respect to $\bar{\mathbf{w}}$ given a fixed $q_{i,l},\forall i,l$ [24]. Thus, although (20b) is a non-convex constraint, its left-hand side consists of a convex function. These properties motivate us to adapt the SCA method [37] to optimize the problem iteratively. Specifically, the constraints can be approximated by their first-order Taylor expansion.
Let us proceed by rewriting $v_{o,m}$ considering fixed $q_{i,l}$. For this, we define $\mathbf{a}_{m,n}=\boldsymbol{\gamma}_{m,n}\odot\mathbf{q}\odot\mathbf{h}$ and leverage the fact that the average power of a signal is equal to the power of its spectrum. Thus, we have [6]
$\mathbb{E}_{t}\bigl{\{}y_{m}(t)^{2}\bigr{\}}=g_{m,1}(\{\bar{\mathbf{w}}_{n}\}_{\forall n})=\frac{G^{2}}{2}\sum_{n}|\mathbf{a}^{T}_{m,n}\bar{\mathbf{w}}_{n}|^{2},$ (21)
$\mathbb{E}_{t}\bigl{\{}y_{m}(t)^{4}\bigr{\}}=g_{m,2}(\{\bar{\mathbf{w}}_{n}\}_{\forall n})=\frac{3G^{4}}{8}\sum_{\begin{subarray}{c}n_{0},n_{1},n_{2},n_{3}\\ n_{0}+n_{1}=n_{2}+n_{3}\end{subarray}}...$
@@
Theorem 1.
The objective function in (20a) can be upper bounded by using
$\sum_{i=1}^{N_{rf}}\mathbb{E}_{t}\biggl{\{}\sqrt{\sum_{l=1}^{N_{h}}x^{DMA}_{i,l}(t)^{2}}\biggr{\}}\leq\sum_{i=1}^{N_{rf}}\sqrt{\sum_{l=1}^{N_{h}}\sum_{n=1}^{N_{f}}\frac{G^{2}}{2}|\omega_{i,n}q_{i,l}h_{i,l}|^{2}},$ (25)
where the right-hand side of the inequality is convex w.r.t. $\omega_{i,n}$.
Proof.
The proof is provided in Appendix -A. ∎
@@
$\operatorname*{minimize}_{\{\bar{\mathbf{w}}_{n}\}_{\forall n}}\quad$ $\Upsilon\!=\!\frac{\sqrt{P_{max}}}{\bar{\eta}}\sum_{i=1}^{N_{rf}}\sqrt{\sum_{l=1}^{N_{h}}\sum_{n=1}^{N_{f}}\frac{G^{2}}{2}|\omega_{i,n}q_{i,l}h_{i,l}|^{2}}+P_{in}$ (26a)
@@
$\sum_{n}\tilde{g}_{m,n}(\bar{\mathbf{w}}_{n}^{(0)})\bigl{(}\bar{\mathbf{w}}_{n}-\bar{\mathbf{w}}_{n}^{(0)}\bigr{)},\quad\forall m,$ (26b)
$\bar{\mathbf{w}}_{n}[(i-1)N_{h}+l]=\bar{\mathbf{w}}_{n}[i],\forall i,l=1,\ldots,N_{h}.$ (26c)
Notice that by utilizing (25) and the fact that (20a) consists of two positive terms, one has that (26a) serves as an upper bound for (20a). Moreover, the inequality in (23) ensures that the solution to this problem is a feasible solution of (20) at each point. Interestingly, the problem has become convex and can be solved at a given point by standard convex optimization tools, e.g., CVX [38]. [^3] For the sake of easy notation and facilitating the reader’s understanding, we have kept the problem in vector form and introduced the constraint (26c) into the problem. However, the problem can be easily converted to scalar form, which removes the mentioned constraint. Moreover, the solution can be iteratively updated using the SCA algorithm [37].